COMMUTING POLYNOMIALS AND

POLYNOMIALS WITH SAME JULIA SET

Pau Atela Jun Hu

Abstract. It has been known since Julia that polynomials commuting under composition have the same Julia set. More recently in the works of Baker and Eremenko, Fernández, and Beardon, results were given on the converse question: When do two polynomials have the same Julia set? We give a complete answer to this question and show the exact relation between the two problems of polynomials with the same Julia set and commuting pairs.

Typeset by AMS-TEX